OpenStudy (melody323):

What type of triangle is shown below? acute triangle right triangle obtuse triangle equiangular triangle

OpenStudy (melody323):

All acute triangles are equiangular. True False

OpenStudy (anonymous):

the angles are not all the same measurements so it isn't d. the triangle isn't made of only acute angles so it isn't a. I think it is c obtuse angle but I am not sure because I do not have a protractor on hand.

OpenStudy (melody323):

Thank You.

OpenStudy (anonymous):

im thinking that not all acute angles are equiangular because it would mean that all three angles of the triangle are 60 degrees and the definition of an acute triangle is anything less than 90

Directrix (directrix):

@jiteshmeghwal9 109 + 52 + 33 does not sum to 180.

OpenStudy (anonymous):

the x would equal 24 @jiteshmeghwal9 so 4(24)+5=101 2(24)=48 (24)+7=31 101+48+31=180 therefore the triangle would be C obtuse triangle

OpenStudy (jiteshmeghwal9):